OpenStudy (anonymous):

Given the function \[x_1^{x_2^{x_3}}\]. Take the derivative of \[x_1, x_2, x_3\] \[f \prime_{x_1}x = x^{x_2x_3} \rightarrow x_2x_3*x_1^{(x_2x_3)-1}\] I don't understand how to take the derivative of \[x_2\] and \[x_3\].. Some guidelines would be appreciated. :D

hartnn (hartnn):

do u know chain rule ?

OpenStudy (anonymous):

\[f(g(x)) = f\prime(g(x))*g\prime x\]

hartnn (hartnn):

yes, [for diff. w.r.t x2] so, x1 and x3 are constants and x2 is variable. for simplicity, i'll take the function as \(a^{x^b}\) using chain rule, it'll be \((a^{x^b})' = (a^{x^b})\ln x^b(x^b)'\) got this ?

hartnn (hartnn):

a =x1, x=x2, b=x3

hartnn (hartnn):

doubts ?

hartnn (hartnn):

its easier for diff. w.r.t x3 x1 and x2 are constants and x3 is variable. for simplicity, i'll take the function as \(a^{b^x} \) let a^b = c a new constant. then you'll just need to differentiate c^x substitute back in c = a^b a=x1,b=x2,c=x3

hartnn (hartnn):

*x=x3

OpenStudy (anonymous):

Been scribbling on my own here and I guess I got it.

hartnn (hartnn):

GOOD :)